OpenStudy (anonymous):

find the logarithm log4(256)

OpenStudy (anonymous):

154.12735778

OpenStudy (anonymous):

Power rule will make it simple.

OpenStudy (anonymous):

\[\log_4(256)\]

OpenStudy (anonymous):

And the power rule says: \[\log(a)^m = m \log(a)\]

OpenStudy (anonymous):

So, you can write \(256\) as \(4^4\)

OpenStudy (anonymous):

\[\log_4(256) = \log_4(4)^4 \implies \color{green}{4} \log_4(4)\]

OpenStudy (anonymous):

and one more thing: \[\log_a(a) = 1\]

OpenStudy (anonymous):

So: \[\log_4(4) = 1\]

OpenStudy (anonymous):

So, at last you have: \[\color{blue}{\log_4(256) = 4}\]